Resistor 1 of 7 Ohms, resistor 2 of 6 Ohms, and resistor 3 of 13 Ohms are connected purely in series to a battery. The current through resistor 1 is 4.9 amps. What is the power dissipated by resistor 2 in Watts?
Answer #1

Equivalent resistance, R = 7 + 6 + 13 = 26 ohm

Power, P = I^2 R = 4.9^2 x 6 = 144.06 W

Comment in case any doubt please rate my answer....